OpenStudy (anonymous):

given that \[f(x)=\ln(2\cos(\frac{x}{2}))\] has a fourier series \[f(x)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1} \cos(nx)}{n}\]find \[\sum_{n=1}^{\infty}\frac{(-1)^n}{n}\]

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=1+-+1%2F2+%2B+1%2F3+-+1%2F4+%2B+1%2F5+-+... Put x=0 and get your result \[ \sum_{n=1}^{\infty}\frac{(-1)^{n+1} \cos(n \times 0 )}{n} = \ln(2\cos(\frac{0}{2}))\]

OpenStudy (anonymous):

OH Thanks... I dont know why my mind is blocked i thought too far of it

OpenStudy (experimentx):

the job of mathematics is to make simple ... most of the problems (very difficult one) i've faced so far had very simple solutions the other way to solve this is to expand ln(1+x) and put x=1

OpenStudy (anonymous):

Yes i know, but this was a question on my test, and it said 10 marks so i couldnt think simple

OpenStudy (experimentx):

lol ... it was more like intimidation ... i certainly find Fourier series quite difficult

OpenStudy (anonymous):

i prefer fourier series in Signals and systems lol not in mathematics